# Enhancing LLMs for Power System Simulations: A Feedback-driven Multi-agent Framework

### v1

@@ V-C Evaluation on MATPOWER
The evaluation results on MATPOWER are illustrated in Figs. 9, 10, and 11. Specifically, Fig. 9 presents the scores achieved by each evaluated scheme in individual attempts when managing complex tasks. Fig. 10 shows the distribution of scores across attempts for each scheme, and Fig. 11 depicts the success rates of each scheme. The outcomes observed here align closely with the results on Daline, enabling us to only focus primarily on comparative analyses across schemes in the subsequent discussion.
@@
TABLE II: Average Cost Analysis of GPT4o-Full Per Task
@@
V-D Cost Analysis
The cost analysis of GPT4o-Full for executing simulation tasks in Daline and MATPOWER is presented in Table II, with average values across all tasks shown. The reported time covers the entire process, including retrieval, reasoning, code generation, simulation execution, result aggregation, and, where necessary, code correction. Remarkably, GPT4o-Full completes each task in approximately half a minute. Additionally, the token expense per task is roughly 0.014 USD.
It is noteworthy that aside from parallel retrieval, no specialized acceleration techniques were employed in this framework. Thus, despite its already satisfactory performance, there is considerable potential for speed enhancements. Even in the current state, without any specific acceleration strategies, GPT4o-Full can execute approximately 120 simulation tasks per hour in Daline and MATPOWER, with a total token cost of around 1.68 USD. Given this high efficiency and cost-effectiveness, our framework presents a promising pathway to improve researcher productivity.
VI Conclusion
This paper addresses the research gap in enhancing LLMs for power system simulations by proposing a feedback-driven, multi-agent framework, representing the first systematic approach to significantly improve LLMs’ simulation capabilities across both familiar and new tools. Validated on 69 diverse simulation tasks from Daline and MATPOWER, our framework achieved substantial performance improvements, with success rates of 93.13% and 96.85%, respectively, far surpassing those of baseline schemes, including the latest LLM, o1-preview. Key findings include: (i) The original simulation capability of LLMs is limited, as evidenced by GPT4o and o1-preview achieving success rates no higher than 27.77%. (ii) Even with the standard RAG module and a comprehensive knowledge base, LLMs achieve overall success rates below 45%, highlighting the need for a more comprehensive approach. (iii) Our framework’s high success rate stems from a synergistic integration of enhanced RAG, enhanced reasoning, as well as environmental acting and feedback mechanisms. Removing any of these elements results in a significant performance decline. (iv) Our framework enables LLMs to execute tasks efficiently, with each task completed in approximately 30 seconds at a token cost of only 0.014 USD, offering a scalable, cost-effective solution that enhances productivity of human scientists in power systems.
However, several critical future challenges still remain: (i) Developing automatic evaluation methods for unbenchmarked results to improve reliability and autonomy. (ii) Expanding the framework to synchronize multiple simulation tools to tackle more challenging tasks. (iii) Since 100% accuracy remains unachieved, integrating error-detection mechanisms to flag uncertainties and inform researchers of potential inaccuracies. Overall, our work is an initial step in the long journey toward realizing intelligent LLM-based research assistants, with potential ahead.

### v2

@@ V-C Evaluation on MATPOWER
The evaluation results on MATPOWER are illustrated in Figs. 9, 10, and 11. Specifically, Fig. 9 presents the scores achieved by each evaluated scheme in individual attempts when managing complex tasks. Fig. 10 shows the distribution of scores across attempts for each scheme, and Fig. 11 depicts the success rates of each scheme. The outcomes observed here align closely with the results on Daline. This alignment allows us to focus primarily on comparative analyses across schemes in the subsequent discussion.
@@
Figure 12: Scores achieved by each evaluated scheme in individual attempts when handling standard tasks (simulation environment: MATPOWER; GPT4o version: gpt-4o-2024-08-06).
V-D Comparison with Supervised Fine-tuning
To further verify the proposed framework’s performance, it was compared with supervised fine-tuning (SFT) [^2] The OpenAI’s SFT approach was used in this paper; see [here] for details., a well‐established approach known to enhance LLM performance on specific tasks—including unseen tasks such as translation and natural language inference [27]. In order to precisely demonstrate the improvements introduced by SFT, GPT4o-Sole has been employed as the foundation for SFT training [^3] Since gpt-4o-2024-08-06 is the only version available for SFT, all GPT4o-related schemes in Section-VD use this version for consistency., denoted by GPT4o-Sole-SFT in Table I.
The initial comparison was conducted on MATPOWER tasks. For this evaluation, 50 random tasks (following the recommendation from OpenAI) were generated with preferred coding responses, producing a supervised training dataset of 24,249 tokens. The model was fine-tuned for 3 epochs, with a batch size of 1 and a learning rate multiplier of 2 (all automatically configured). Fig. 12 illustrates the performance of GPT4o-Full, GPT4o-Sole, and GPT4o-Sole-SFT, based on the scores achieved in individual attempts on standard tasks. Evidently, SFT enhances the performance of GPT4o-Sole: GPT4o-Sole-SFT consistently attains more perfect scores (100 points) than GPT4o-Sole, with success rates of 51.11% versus 35.18%, respectively. However, GPT4o-Full significantly outperforms GPT4o-Sole-SFT, achieving 100 points for all tasks starting from the second attempt.
The underperformance of GPT4o-Sole-SFT could stem from the representativeness of the training datasets—despite significant human effort in generating them, these datasets may not and cannot capture every nuanced detail of MATPOWER. To rigorously eliminate the possibility that the limited dataset is responsible for the performance gap, an additional evaluation was conducted using the Daline simulation environment. In this experiment, the full testing dataset (i.e., the Daline simulation tasks used for evaluation) was used as the SFT training dataset. Specifically, all 34 testing simulation tasks, augmented with 16 additional tasks, were employed for fine-tuning, yielding a training dataset of 32,346 tokens. GPT4o-Sole-SFT was fine-tuned for 3 epochs, with a batch size of 1 and a learning rate multiplier of 2 (again, all automatically configured). The score distributions and specific success rates are presented in Fig. 13 and Table III. This evaluation shows that GPT4o-Sole-SFT improves the success rate of GPT4o-Sole from 0% to 51.96% overall, with 28.57% for complex tasks and 58.03% for standard tasks. Yet, even with the entire testing dataset used for fine tuning, GPT4o-Sole-SFT still fails to reach high accuracy.
The underperformance of GPT4o-Sole-SFT can be attributed to the inherent limitations of SFT for LLMs: SFT can capture coarse‐grained information such as style, tone, format, or other qualitative aspects, but SFT cannot memorize every nuance of the training data. The reason is twofold: (1) SFT is a lossy compression process, which inevitably discards finer details, and (2) SFT only adjusts a small subset of the model’s parameters, with the rest remaining focused on generating generalizable patterns — as a result, SFT cannot achieve 100% retention of task-specific details, especially for tasks like coding, which require high-precision parameters and functions. In contrast, GPT4o-Full leverages an enhanced RAG module, allowing the model to dynamically access and retrieve precise information from external documents. This retrieval mechanism effectively mitigates the limitations of parameter-based memorization, enabling GPT4o-Full to achieve over 95% accuracy across both complex and standard tasks.
Figure 13: Distribution of scores achieved across attempts for each evaluated scheme (simulation environment: Daline; GPT4o version: gpt-4o-2024-08-06).
TABLE III: Success rates for each scheme, broken down by all tasks combined, complex tasks only, standard tasks only, as well as for the first attempt success rate and the final attempt success rate (simulation environment: Daline; GPT4o version: gpt-4o-2024-08-06)
Schemes GPT4o-Sole GPT4o-Sole-SFT GPT4o-Full
All-Tasks 0.000% 51.961% 95.098%
All-Complex 0.000% 28.571% 95.238%
All-Standard 0.000% 58.025% 95.062%
First Attempt 0.000% 50.000% 91.176%
Final Attempt 0.000% 52.941% 97.059%
V-E Cost Analysis
The cost analysis of GPT4o-Full for executing simulation tasks in Daline and MATPOWER is presented in Table IV, with average values across all tasks shown. The reported time covers the entire process, including retrieval, reasoning, code generation, simulation execution, result aggregation, and, where necessary, code correction. Remarkably, GPT4o-Full completes each task in approximately half a minute. Additionally, the token expense per task is roughly 0.014 USD.
It is noteworthy that aside from parallel retrieval, no specialized acceleration techniques were employed in this framework. Thus, despite its already satisfactory performance, there is considerable potential for speed enhancements. Even in the current state, without any specific acceleration strategies, GPT4o-Full can execute approximately 120 simulation tasks per hour in Daline and MATPOWER, with a total token cost of around 1.68 USD. Given this high efficiency and cost-effectiveness, the proposed framework presents a promising pathway to improve researcher productivity.
It must be emphasized, however, that the reported execution time and token cost serve only as approximate indicators, given the inherent variability introduced by factors such as retrieval, reasoning, code execution, result aggregation, and iterative error correction. This analysis aims to provide readers with a general sense of the efficiency of automated simulations within the scope of the specific simulation tasks, rather than to establish precise quantitative relationships.
VI Conclusion
This paper addresses the research gap in enhancing LLMs for power system simulations by proposing a feedback-driven, multi-agent framework. It represents the first systematic approach to significantly improve LLMs’ simulation capabilities across both familiar and new tools. Validated on 69 diverse simulation tasks from Daline and MATPOWER, the proposed framework achieved substantial performance improvements, with success rates of 93.13% and 96.85%, respectively. It far surpasses those of baseline schemes, including the latest LLM, o1-preview, and SFT for LLMs. Key findings include: (i) The original simulation capability of LLMs is limited, as evidenced by GPT4o and o1-preview achieving success rates no higher than 27.77%. (ii) Even with the standard RAG module and a comprehensive knowledge base, LLMs achieve overall success rates below 45%, highlighting the need for a more comprehensive approach. (iii) Our framework’s high success rate stems from a synergistic integration of enhanced RAG, enhanced reasoning, as well as environmental acting and feedback mechanisms. Removing any of these elements results in a significant performance decline. (iv) While SFT may be able to capture coarse-grained attributes such as style, tone, and format, it struggles to retain the full detail necessary for power system simulations. Even when the entire test dataset is included in the training set, SFT achieves a success rate of less than 60%. This limitation stems from its inherently lossy compression mechanism and constrained parameter update capacity, which prevent it from encoding fine-grained details with high fidelity. (iv) The proposed framework enables LLMs to execute tasks efficiently, with each task completed in approximately 30 seconds at a token cost of only 0.014 USD (for the simulation tasks used in this paper), offering a scalable, cost-effective solution that enhances productivity of human scientists in power systems.
TABLE IV: Average cost analysis of GPT4o-Full per task
@@ VI Conclusion
  • $\dagger$
    A token is a fundamental unit of text, typically representing fragments of words. Tokens quantify both the input (text submitted to the LLM via API) and the output (text or code generated by the LLM via API).
  • $\ddagger$
    Expense refers to the monetary cost incurred for processing these tokens, as charged by the API provider (OpenAI in this study).
    The API provider automatically computes the total tokens and associated expenses. The average values here are calculated by dividing the total tokens (and corresponding expenses) used across all simulation tasks with GPT4o-Full by the number of tasks.
However, several critical future challenges still remain. First, while the proposed framework demonstrates significant improvements over existing schemes and supervised fine-tuning, achieving a perfect 100% success rate for LLM-based simulations remains an open challenge. For instance, the proposed framework sometimes fails to detect “non-execution-bug” failures, resulting in unaware inaccurate outputs that cannot be automatically corrected. A potential avenue for future work is to explore the use of a RAG-supported, fine-tuned LLM-based code-checking agent, which would verify executed simulation code and identify hidden errors that the error-reporting mechanism may overlook. Second, when simulation requests are inherently ambiguous or underspecified, accurately determining the user’s intent becomes difficult for any automated method—not just the framework presented in this paper. This ambiguity primarily arises from the initial lack of detail in the user’s input, rather than any intrinsic limitation of natural language coding. To address this, a promising future direction is to integrate an interactive dialogue stage into the framework. In this approach, an autonomous agent would evaluate the specificity of the initial request and, if necessary, prompt the user with targeted clarifying questions. This human-in-the-loop strategy could help refine the requirements, leading to more precise retrieval and improved code generation.